# Epistemic diversity across language models mitigates knowledge collapse

### v1

@@ Abstract.
The growing use of artificial intelligence (AI) raises concerns of knowledge collapse, i.e., a reduction to the most dominant and central set of ideas. Prior work has demonstrated single-model collapse, defined as performance decay in an AI model trained on its own output. Inspired by ecology, we ask whether AI ecosystem diversity, that is, diversity among models, can mitigate such a collapse. We build on the single-model approach but focus on ecosystems of models trained on their collective output. To study the effect of diversity on model performance, we segment the training data across language models and evaluate the resulting ecosystems over ten, self-training iterations. We find that increased epistemic diversity mitigates collapse, but, interestingly, only up to an optimal level. Our results suggest that an ecosystem containing only a few diverse models fails to express the rich mixture of the full, true distribution, resulting in rapid performance decay. Yet distributing the data across too many models reduces each model’s approximation capacity on the true distribution, leading to poor performance already in the first iteration step. In the context of AI monoculture, our results suggest the need to monitor diversity across AI systems and to develop policies that incentivize more domain- and community-specific models.
artificial intelligence, epistemic diversity, model collapse, AI ecosystem, system pluralism, distributed artificial intelligence
@@ 1. Introduction
Figure 1. Illustration of model collapse in AI ecosystems. Because of a model’s (m) inherent biases (in learning and resampling) and finite data, it cannot perfectly regenerate the information space (e.g., concepts, claims, facts) it was trained on. Some information becomes distorted, and some is lost entirely due to low sampling probabilities. When this training–generation cycle is repeated, the model’s output becomes increasingly homogeneous and degrades over iterations $t$. The probability distribution of the information space expressed in the output (and represented by the model, $p_{\theta}$) gradually narrows with increasing $t$, converging toward delta-like functions. Collapse can occur at both the individual model level and the ecosystem level, meaning that even a collection of models trained on their collective output can eventually degenerate (here illustrated with three models). In society, we expect knowledge to grow and improve. In contrast, without human intervention, the best we can hope for from AI models is that they preserve the distribution of information they are originally trained on ($p_{data}$). Thus, for the purpose of this paper, we evaluate ecosystems based on model performance on the original dataset, averaged over 10 training–generation iterations.
@@
AI’s growing role in knowledge production and curation raises concerns about a potential collapse of human knowledge (Peterson, 2025; Sourati et al., 2025; Wagner and Jiang, 2025). The scenario unfolds as follows: as AI systems become widely incorporated in private and professional contexts (e.g., in scientific discovery and writing (Liang et al., 2025; Mishra et al., 2024; Messeri and Crockett, 2024)), they also generate an increasing volume of artificial data used for training. This amplifies the risk of model collapse, a degenerative process in which generative models re-trained on their own outputs begin producing homogeneous, biased, and nonsensical outputs (Herel and Mikolov, 2024; Shumailov et al., 2024; Bertrand et al., 2024). Ultimately, such homogenization in model-generated data, particularly with language, may contribute to knowledge collapse, a reduction of human knowledge to the most dominant and central set of ideas (Peterson, 2025; Wright et al., 2025).
@@
Even before AI, philosophers have argued for the benefits of epistemic diversity in science, not only for advancing knowledge production but also to prevent injustices such as the undervaluing of minority perspectives (Douglas, 2009; Longino, 1990; Wright et al., 2025; Rolin, 2019; Solomon, 2006). When scientists learn from both their own results and one another’s findings, diversity reduces the risk of converging on false beliefs. For example, the myth that Sickle Cell Disease is specific to Black Americans persisted in medicine for decades (Anderson et al., 2023). Simulations of collective knowledge production show that diversity is beneficial, not monotonically, but up to a certain level; if diversity increases to the point where scientists no longer learn from each other, the overall success rate begins to decline as well [^1] To clarify our position: We do not believe this upper bound has ever been reached. (Zollman, 2010; Wu and O’Connor, 2023; Fazelpour and Steel, 2022).
@@
Like scientists exchanging new findings, AI models learn from their own output and the output of other AI [^2] We are not attempting to compare humans and AI models underlying learning mechanisms but only to illustrate information interactions in human and AI ecosystems. We use terms like “learning” and “training” because they are established in AI research.. In the presence of AI monoculture—where a few large models, trained on similar datasets with similar model architectures, (Lu and Kleek, 2024; Lazar and Nelson, 2023; Sourati et al., 2025; Bommasani et al., 2022) dominate the AI ecosystem—we ask whether such an optimal degree of diversity exists in AI ecosystems that could mitigate monocultural shifts.
@@
In homogeneous ecosystems, model collapse occurs because of a model’s non-zero probability of losing information or learning wrong patterns, which is irreversible, when relying only on its own output (Shumailov et al., 2024). Analogous to diversity in human systems, we hypothesize that diverse models trained on their collective outputs can benefit their performance by increasing distributional variance in artificial data generation. Viewing an AI model as an epistemological representation, we use the Hill-Shannon Diversity (HSD, $D$) (Roswell et al., 2021; Jost, 2006; Hill, 1973) to measure ecosystem diversity, which quantifies the effective number of equally frequent, diverse AI models. It is important to note that we manipulate only the diversity across models, without altering the diversity of the dataset on which the models are initially trained, as has been studied in prior work (e.g., Wright et al., 2025; Schaffelder and Gatt, 2025; Zhu et al., 2025).
@@
To address the challenges of studying the evolution of the entire AI ecosystem (e.g., ChatGPT, Gemini, Claude, etc., and all their inputs and outputs), we build upon previous work that analyzes a single-model collapse (e.g. Shumailov et al., 2024; Herel and Mikolov, 2024). This previous work re-trains small language models on their own output over multiple iterations and measures performance using perplexity. We take a similar approach, but instead focus on a collection of models, or a model ecosystem, rather than a single model. This then allows us to study the effects of our main variable of interest—ecosystem diversity—on model performance. We measure this effect separately on two different model types (OPT-125m (Zhang et al., 2022) and GPT2 (Radford et al., 2019)).
@@
While there are many forms of diversity across models (e.g., size, architecture, alignment, system messaging) [^3] A specific model architecture can be used to build different AI systems (i.e., the end products with which users interact). For simplicity, we use “model type” to differentiate between OPT-125m and GPT2 pre-trained base models and “model” to refer to the fine-tuned versions., we focus on training data as the way of instantiating epistemic diversity. To do this under a controlled setting, we segment the training data, which is fixed in size and, in the first iteration step, also in content. Through this process, we go from one, pre-trained model to multiple fine-tuned models (MM=1, 2, 4, 16), as shown in Figure 2. We use Wikitext2 (Merity et al., 2016) as the dataset for generating artificial training data, evaluation, and initial training. This segmentation approach allows us to vary ecosystem diversity ($D=M$) in a relatively simple manner, while keeping factors, such as model type and the per-ecosystem resource budget for sample size and computation, fixed.
Our experiments under a fixed budget yield an optimal ecosystem diversity greater than zero for model performance when models are trained on their own output, as shown in Figure 3. Too much diversity (and less training data per model) results in high perplexity at early stages, and too little leads to a high degeneration rate in the long term. We theorize that a trade-off primarily between increasing ecosystem-level expressivity (the ensemble can represent a richer mixture as members can specialize to subpopulations) and decreasing approximation capacity of the individual training data on the true distribution explains the U-shaped relationship between performance and diversity.
@@
The optimal ecosystem diversity observed in our experiments suggests that a similar optimum may exist in reality. As AI models become increasingly incorporated into major institutions such as science, education, government, and economics, a collapse toward the most dominant ideas could have severe societal implications. These risks extend beyond knowledge production and may also impact domains of epistemic and social justice.
@@ 2.1. Collapse
Model collapse has been theoretically and/or empirically demonstrated for generative models in various settings (Alemohammad et al., 2023; Shumailov et al., 2024; Herel and Mikolov, 2024; Zhang et al., 2025b; Xing et al., 2025; Dohmatob et al., 2024b). The observed effects typically involve some form of performance decay; however, the field has not yet arrived at a general definition of the phenomenon (Schaeffer et al., 2025). Our experimental approach for the single model follows Shumailov et al. (2024), who first formalized model collapse and revealed it in OPT-125m. Model collapse arises primarily from three sources of error related to training data quality (precision and recall), the model’s capacity to represent a given distribution (to “express” it), and the model’s capacity to accurately learn what it is capable of representing (Shumailov et al., 2024): statistical approximation error, functional expressivity error, and functional approximation error. We will return to these errors when discussing differences in model performance across collections of diverse models. Model collapse is not only a threat to generative models accidentally (or inevitably) trained on artificial internet data but also when considering approaches leveraging synthetic data for efficiency reasons (Bai et al., 2022; Goyal and Mahmoud, 2024; Feng et al., 2024a). Very recent work theorizes that model collapse will result in a collapse of knowledge, defined as the continuous narrowing down of ideas, which will lead to broader epistemic problems for our society (Wright et al., 2025; Peterson, 2025; Wagner and Jiang, 2025).
2.2. Mitigation and prevention
Because of the epistemic risks of collapse, scholars have developed various strategies to prevent such performance decay, or, put differently, to ensure stability.
The majority of intervention strategies focus on data curation, primarily at the data level, including adding fresh data and filtering or correcting artificial data (Gerstgrasser et al., 2024; Bertrand et al., 2024; Amin et al., 2025; Xing et al., 2025; Fu et al., 2025; Zhu et al., 2025; Dohmatob et al., 2024b; Feng et al., 2024a). While such methods have been shown to mitigate the problem, in most experimental settings they cannot entirely prevent performance decay, even when the involved technical and practical challenges of constantly sourcing fresh data or reliably separating real from artificial data set aside (Xing et al., 2025; Dohmatob et al., 2024a).
@@ 2.3. AI ecosystem
Our real AI ecosystem does not consist of a single self-training model but rather a collection of models trained on both their own outputs and those produced by others. Thus, more closely aligned with the work presented here are recent approaches that seek to understand model collapse at the ecosystem level––for instance, by simulating interactions among AI models via retrieval-augmented generation (RAG) (Wang et al., 2025), or by sourcing artificial data from multiple language models for training and evaluating a single model (Schaffelder and Gatt, 2025). In contrast, our approach focuses on diversity across models as the independent variable, involving real (non-simulated) re-training of language models on their collective output.
@@ 2.4. Diversity
AI models generate information according to learned probability distributions, which means each model reflects specific assumptions about what counts as relevant, likely, or meaningful knowledge; in other words, a particular epistemology. Epistemic diversity in an AI ecosystem therefore captures the extent to which multiple models, shaped by different data sources, architectures, or objectives, may yield divergent interpretations or outputs when presented with the same input. In recent AI research, epistemic diversity within and across AI models has attracted increased interest. For example, scholars have shown that diverse models collectively make better decisions (Kleinberg and Raghavan, 2021), reduce the risk of failures (Lu and Kleek, 2024), and enable fairer representations of diverse societies (Fisher et al., 2025; Varon et al., 2024; Gebru and Torres, 2024; Jain et al., 2024; Blili-Hamelin et al., 2025).
Originating in ecology, a common method to measure such diversity is the Hill–Shannon Diversity (HSD) (Roswell et al., 2021; Jost, 2006; Hill, 1973; Wright et al., 2025), which reflects the effective number of diverse, equally common elements (e.g., models, species, or categories) in a population.
Figure 2. Instantiating ecosystem diversity and evaluating model performance. We examine the performance of models $m$ within collections of increasing diversity $D$ when trained on collective output across 10 training iterations $t$. The number of models $M$ equals the Hill-Shannon Diversity $D$, which is instantiated through segmentation of the collected output. At each iteration, we concatenate and shuffle the $M$ generated datasets, and uniformly redistribute to the $M$ models for the next step. This segmentation approach keeps the number of tokens per ecosystem NN fixed, yielding a per-model sample size of $n=N/M$. The ground-truth dataset used to measure performance, the original Wikitext2 test split, is fixed across all models and iterations.

### v2

@@ Abstract.
As artificial intelligence (AI) becomes more widely used, concerns are growing that model collapse could lead to knowledge collapse, i.e. a degradation to a narrow and inaccurate set of ideas. Prior work has demonstrated single-model collapse, defined as performance decay in an AI model trained on its own outputs. Inspired by ecology, we ask whether increasing AI ecosystem diversity (i.e., the number of distinct models) can mitigate such collapse. To study the effect of diversity on model performance, we extend the single-model approach by segmenting the training data across an increasing number of language models and evaluating the resulting ecosystems of models over ten self-training iterations. We find that training a single model on the entire dataset improves performance only in the short term but amplifies collapse over longer horizons. Specifically, we observe that the optimal diversity level (i.e., the level that maximizes performance) increases monotonically with the number of self-training iterations. The observed effect is robust across various experimental settings, including different model families, parameter sizes, mixing human- and model-generated data, and temperature sampling methods, demonstrating the significance of ecosystem diversity for mitigating collapse. Moreover, our experiments with increased model and dataset sizes indicate that scaling up the system can amplify collapse in highly homogeneous ecosystems, thereby increasing the diversity benefits. In light of AI monoculture concerns, our results suggest the need to monitor (dis)agreement among AI systems and to incentivize more domain- and community-specific models to ensure successful knowledge production in the long run.
artificial intelligence, epistemic diversity, knowledge collapse, model collapse, AI ecosystem, system pluralism, AI pluralism
@@ 1. Introduction
Figure 1. Illustration of model collapse in AI ecosystems. Because of a model’s (m) inherent biases (in learning and resampling) and finite data, it cannot perfectly regenerate the information space (e.g., concepts, claims, facts) it was trained on. Some information becomes distorted, and some is lost entirely due to low sampling probabilities. When this training–generation cycle is repeated, the model’s output becomes increasingly homogeneous and degrades over iterations $t$. The probability distribution of the information space expressed in the output (and represented by the model, $p_{\theta}$) gradually narrows with increasing $t$, converging toward delta-like functions. Collapse can occur at both the individual model level and the ecosystem level, meaning that even a collection of models trained on their collective output can eventually degenerate (here illustrated with three models). In society, we expect to advance knowledge. Without human intervention, the best we can hope for from AI models, though, is that they preserve the original distribution of information. Thus, for the purpose of this paper, we evaluate ecosystems based on model performance on a fixed ground-truth dataset over 10 self-training iterations.
@@
The growing role of Artificial Intelligence (AI) in knowledge dissemination and curation raises concerns about a collapse in human knowledge (Peterson, 2025; Sourati et al., 2025; Wagner and Jiang, 2025). The scenario unfolds as follows (Figure 1): as AI systems become widely incorporated in private and professional contexts (e.g., in scientific discovery and writing (Liang et al., 2025; Hao et al., 2026; Messeri and Crockett, 2024)), they also generate an increasing volume of artificial data used for training. This amplifies the risk of model collapse, a degenerative process in which generative models re-trained on their own outputs begin producing homogeneous, biased, and nonsensical outputs (Herel and Mikolov, 2024; Shumailov et al., 2024; Bertrand et al., 2024). Ultimately, such homogenization in model-generated data, particularly with language, may contribute to knowledge collapse, a degradation to a narrow, and potentially inaccurate, set of ideas (Peterson, 2025; Wright et al., 2025; Messeri and Crockett, 2024).
@@
Independent of these recent AI-related concerns, philosophers have argued for the benefits of epistemic diversity in science, not only for advancing knowledge production, but also to prevent injustices such as the undervaluing of minority perspectives (Douglas, 2009; Longino, 1990; Fricker, 2007; Rolin, 2019; Solomon, 2006). When scientists learn from both their own results and one another’s findings, diversity reduces the risk of converging on false beliefs. For example, monoculture thinking in medicine prolonged the myth that sickle cell disease was limited to Black Americans, leading not only to discrimination against the Black community but also to reduced treatment options for anyone affected by the disease. (Anderson et al., 2023).
“If diversity is good, how much diversity is best?” (Solomon, 2006). Simulations of collective knowledge production identify an upper bound on diversity: once diversity increases to the point where scientists no longer learn from one another, overall success rates begin to decline [^1] To clarify our position: We do not believe this upper bound has ever been reached. (Zollman, 2010; Wu and O’Connor, 2023; Fazelpour and Steel, 2022; Pöyhönen, 2017).
@@
Like scientists exchanging new findings, AI models learn from their own output and the output of other AI [^2] We are not attempting to compare humans and AI models underlying learning mechanisms but only to illustrate information interactions in human and AI ecosystems. We use terms like “learning” and “training” because they are established in AI research but recognize the problems in anthropomorphizing this technology.. In the presence of AI monoculture—where a few large models, trained on similar datasets with similar model architectures, dominate the AI ecosystem (Lu and Kleek, 2024; Lazar and Nelson, 2023; Sourati et al., 2025; Bommasani et al., 2022; Zhang et al., 2025a)—and where evidence of model collapse exists, we ask whether increasing ecosystem diversity (i.e., the number of dissimilar models) could mitigate monocultural shifts, and whether this happens monotonically or to an optimal degree.
@@
In a homogeneous ecosystem, collapse occurs because of a model’s non-zero probability of losing information or learning wrong patterns, which is irreversible, when relying only on its own output (Shumailov et al., 2024). Analogous to productive disagreement among humans, we hypothesize that dissimilar models trained on their collective outputs can improve collective performance by correcting each other’s errors. Viewing an AI model as an epistemological representation, we use the Hill-Shannon Diversity (HSD, $D$) from biology (Roswell et al., 2021; Jost, 2006; Hill, 1973) to measure ecosystem diversity, which quantifies the effective number of distinct AI models. It is important to note that the diversity across models serves as the primary independent variable that we manipulate to examine its effect on performance. This stands in contrast to prior work, which has focused on diversity in data (e.g., Schaffelder and Gatt, 2025; Zhu et al., 2025) and/or has treated diversity as an outcome variable (e.g., Wright et al., 2025).
@@
Given the enormous resources required to train and operate prevailing models (e.g., ChatGPT, Gemini, Claude, etc.), studying the impact of diversity on AI ecosystem evolution at real-world scale is infeasible. And, even if we had infinite resources, many confounding factors could shape model performance, including model and data size (Kaplan et al., 2020; Bertrand et al., 2024), model architecture, and system design. To address these challenges, we build upon previous work that analyzes a simple, single-model in a semi-controlled environment (e.g. Shumailov et al., 2024; Herel and Mikolov, 2024). This previous work re-trains small language models on their own output over multiple iterations and measures performance using perplexity. We take a similar approach, but instead focus on ecosystems of models (i.e., collections of models), rather than a single model. This allows us to study the effects of ecosystem diversity on model performance, which we measure separately on two distinct model families (OPT (Zhang et al., 2022) and GPT2 (Radford et al., 2019)).
@@
While models can be differentiated in many ways, (e.g., by size, architecture, alignment, system messaging) [^3] A specific model architecture can be used to build different AI systems (i.e., the end products with which users interact) (Brown et al., 2005). For simplicity, we use “model family” to differentiate between OPT and GPT2 pre-trained base models and “model” to refer to the fine-tuned versions., we instantiate ecosystem diversity through variation in training data.. To do this under a controlled setting, we segment the training data, which is fixed in size and, in the first iteration step, also in content. Through this process, we go from one, pre-trained model to multiple fine-tuned models (MM=1, 2, 4, 16), as shown in Figure 2. We use Wikitext2 (Merity et al., 2016) as the dataset for generating artificial training data, evaluation, and initial training. This segmentation approach allows us to vary ecosystem diversity ($D=M$) in a relatively simple manner, while keeping factors, such as model family and the per-ecosystem resource budget for sample size and computation, fixed. In our primary experimental setting, we use the smallest versions of the two model families (OPT-125m and GPT2).
Our experiments identify an optimal level of ecosystem diversity for model performance, which increases monotonically with the number of self-training iterations (Figure 3). In other words, the prevailing strategy of training a single model on as much data as possible is short-sighted: over longer horizons, a collection of models trained on segmented datasets outperforms a single model. As the number of self-training loops increases, stronger segmentation is required to reach the diversity level that maximizes overall performance. Put simply, the benefits of diversity grow with the number of iterations.
To assess the effect of ecosystem diversity under more realistic conditions, we introduce two variations of our primary setting. In the first variation (V1), we increase the size of system (i.e., model and data size). Scaling up leads to a stronger collapse in more homogeneous ecosystems, while collapse is attenuated in more diverse settings, resulting in increased diversity benefits. The second variation (V2) enhances the training data by (1) incorporating 10% fresh, original Wikitext2 data at each iteration and (2) by applying temperature sampling with $\tau=0.5,1,2$. Prior work suggests that both real-data incorporation (Shumailov et al., 2024; Gerstgrasser et al., 2024; Xing et al., 2025) and data diversification (Wright et al., 2025; Schaffelder and Gatt, 2025; Zhu et al., 2025) can mitigate collapse. While our results confirm these findings, they also show that the effect of ecosystem diversity outweighs the effects of data quality manipulation. Overall, the experimental variations suggest that the benefits of diversity for long-term performance may persist under real-world AI ecosystem conditions, where models such as ChatGPT are substantially larger, trained on far more data, can be optimized for diverse outputs, and where human-generated data is not replaced but instead accumulates alongside artificial data.
@@
As AI models become increasingly incorporated into major institutions such as science, education, government, and economics, a collapse toward the most dominant ideas could have severe societal implications. The results of our experiments suggest that differentiation of models could be a significant factor in mitigating such collapse.
@@ 2.1. Collapse
Model collapse has been theoretically and/or empirically demonstrated for generative models in various settings (Alemohammad et al., 2023; Shumailov et al., 2024; Herel and Mikolov, 2024; Zhang et al., 2025b; Xing et al., 2025; Dohmatob et al., 2024b). The observed effects typically involve some form of performance decay; however, the field has not yet arrived at a general definition of the phenomenon (Schaeffer et al., 2025). Our experimental approach for the single model follows Shumailov et al. (2024), who first formalized model collapse and revealed it in OPT-125m. Model collapse arises primarily from three sources of error related to training data quality, the model’s capacity to represent and accurately learn a given distribution (Shumailov et al., 2024): statistical approximation error, functional expressivity error, and functional approximation error. Prior work show that increasing model and data sizes can amplify collapse (Wright et al., 2025; Herel and Mikolov, 2024; Dohmatob et al., 2024a, b; Feng et al., 2024a), thereby contrasting established scaling laws in machine learning. Model collapse is not only a threat to generative models accidentally (or inevitably) trained on artificial internet data but also when considering approaches leveraging synthetic data for efficiency reasons (Bai et al., 2022; Goyal and Mahmoud, 2024; Feng et al., 2024a). Very recent work theorizes that model collapse will result in a collapse of knowledge, defined as the continuous narrowing down of ideas, which will lead to broader epistemic problems for our society (Wright et al., 2025; Peterson, 2025; Wagner and Jiang, 2025).
2.2. Collapse mitigation
Because of the epistemic risks of collapse, scholars have developed various strategies to prevent such performance decay, or, put differently, to ensure stability. The majority of intervention strategies focus on increasing data quality. This usually involves adding fresh non–model-generated data (e.g., Bertrand et al., 2024; Gerstgrasser et al., 2024; Fu et al., 2025), filtering or correcting artificial data (e.g., Amin et al., 2025; Dohmatob et al., 2024b; Feng et al., 2024a), or diversifying training data (Zhu et al., 2025; Wright et al., 2025; Schaffelder and Gatt, 2025). While such methods have been shown to mitigate the problem, in most experimental settings they cannot entirely prevent performance decay. In addition, it is often infeasible technically and practically to constantly source with fresh data or reliably separate real from artificial data (Xing et al., 2025; Dohmatob et al., 2024a).
@@ 2.3. AI ecosystem
Our real AI ecosystem does not consist of a single self-training model but rather a collection of models trained on both their own outputs and those produced by others. Thus, more closely aligned with the work presented here are recent approaches that seek to understand model collapse at the ecosystem level—for instance, by simulating interactions among AI models via retrieval-augmented generation (RAG) (Wang et al., 2025a), or by sourcing artificial data from multiple language models for training and evaluating a single model (Schaffelder and Gatt, 2025). In contrast, our approach focuses on diversity across models as the independent variable, involving real (non-simulated) re-training of language models on their collective output.
@@ 2.4. Diversity
AI models generate information according to learned probability distributions, which means each model reflects specific assumptions about what counts as relevant, likely, or meaningful knowledge; in other words, a particular epistemology. Epistemic diversity in an AI ecosystem therefore captures the extent to which multiple models, shaped by distinct data sources, architectures, or objectives, may yield divergent outputs when presented with the same input. In recent AI research, epistemic diversity within and across AI models has attracted increased interest. For example, scholars have shown that diverse models collectively make better decisions (Kleinberg and Raghavan, 2021), reduce the risk of failures (Lu and Kleek, 2024), and enable fairer representations of diverse societies (Fisher et al., 2025; Varon et al., 2024; Gebru and Torres, 2024; Jain et al., 2024; Blili-Hamelin et al., 2025).
Originating in ecology, a common method to measure such diversity is the Hill–Shannon Diversity (HSD) (Roswell et al., 2021; Jost, 2006; Hill, 1973; Wright et al., 2025), which reflects the effective number of distinct, equally common elements (e.g., models, species, or categories) in a population.
Figure 2. Experimental design: manipulating ecosystem diversity and evaluating performance across self-training iterations. We examine the performance of models $m$ within collections of increasing diversity $D$ when trained on collective output across 10 self-training iterations $t$. The number of models $M$ equals the Hill-Shannon Diversity $D$, which is manipulated through segmentation of the training data. At each iteration, we concatenate and shuffle the $M$ generated datasets, and uniformly redistribute to the $M$ models for the next step. This segmentation approach keeps the data size (i.e., number of tokens) per ecosystem NN fixed, yielding a per-model sample size of $n=N/M$. The ground-truth dataset used to measure performance, the original Wikitext2 test split, is fixed across all models and iterations.